HALOGEN BOND AND INTERNAL DYNAMICS IN CClF_{3}-H2O

L. EVANGELISTI, G. FENG and \underline{W. CAMINATI}, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy; P. ECIJA, E.J. COCINERO and F. CAS-TANO, Departamento de Quimica Fisica, Facultad de Ciencia y Tecnologia, Universidad del Pais Vasco (UPV-EHU), Apartado 644, E-48080 Bilbao, Spain.

The pure rotational spectra of CF_{3}^{35}Cl-H_{2}O, CF_{3}^{37}Cl-H_{2}O, CF_{3}^{35}Cl-H_{2}^{18}O, CF_{3}^{35}Cl-OHD, CF_{3}^{35}Cl-D_{2}O isotopologues have been investigated by pulsed supersonic-jet FT-microwave spectroscopy. We assigned the m = 0 and m = 1 (for the first 2 isotopologues) states and the spectra, for all, are characteristic of a symmetric top of evenly spaced band. A substitution analysis was made for the m = 0 state of the dimers with H_{2}^{18}O and D_{2}O. Ab initio calculation (MP2 level of electron correlation and 6-311++G** basis set) have been carried out in order to obtain information about the structure and relative stability. The interaction between the subunits occurs via C-Cl⋯O(H_{2}O) halogen bond.